OpenStudy (anonymous):

SAT subject test question: A triangle has sides measuring 3,5,and 7 inches. What is the measure of its largest angle?

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

The largest angle will be opposite the longest side length. So we can draw it like this.And from here, I think we can use the `Law of Cosines` to solve for angle C.

zepdrix (zepdrix):

\[\large c^2=a^2+b^2-2ab \cos C\]

zepdrix (zepdrix):

Understand how to solve it from here? :D

OpenStudy (anonymous):

yes! Thank you :)